£250 is invested in a saving account. The nominal rate convertible monthly for the first 3 months is 18% and the nominal rate of interest convertible quarterly for the next 9 months is 20%. How much is in the account at the end of the year?
1
Expert's answer
2020-09-29T19:04:36-0400

solution


For the first 3 months:


"Annual\ rate =18\%"

The rate convertible monthly is:


"\frac{18 \%}{12} = 0.015"

The number of compounding periods in the 3 months is 3


For the next 9 months:


"Annual\ rate =20\%"

The rate convertible quarterly is:


"\frac{20 \%}{4} = 0.05"

The number of compounding periods in the 9 months is:


"\frac{9}{12}*4=3"

Therefore the investment grows by 0.015 for 3 periods and then by 0.05 for a further 3 periods.



"Future \ value = P*(1+r_1)^{n1}*(1+r_2)^{n2}"

"= 250*(1.015)^{3}*(1.05)^{3}=302.6259"

answer: the investment will be worth £ 302.63